# Quantum-enhanced Markov Chain Monte Carlo for Combinatorial Optimization

### v1

@@ V Discussion & Conclusions
As quantum computing hardware continues to improve, the reliable execution of quantum optimization experiments for non-trivial systems comes within reach. More explicitly, we can execute experiments that study larger and harder problems, and provide an example pathway for seeking quantum advantage in optimization. This is well-supported by benchmarking which enables the community to explore the strengths of classical and quantum algorithms in solving combinatorial optimization problems. Open-source quantum optimization benchmarking frameworks like QOBLIB, from which the majority of problem instances studied here are sourced, help us to understand how we move from utility scale quantum experiments towards advantage in optimization.
In this manuscript, we explore a quantum solver based on a QeMCMC sampling routine to solve non-trivial combinatorial optimization problem instances. We have adapted the QeMCMC workflow to incorporate a warm-started optimization procedure and intentional bias to our proposal distribution towards “good” solutions. We also combine a parallel tempering approach to explore the available cost landscape more efficiently, which is regularly used to find ground states of many-body systems. This heuristic quantum-classical algorithm has been tested against problem instances with up to 117 decision variables, corresponding to 117 qubits on real quantum hardware. While our method does not have provable convergence guarantees, we recover the global optimum for problem instances with 17, 52, and 117 variables.
The results presented indicate early evidence of a scaling advantage compared to classical MCMC approaches for the chosen problem instances. Interestingly, we also see the quantum hardware outperform MPS simulations for the considered setting. We intend to explore larger problem instances to improve our understanding of the scaling of these methods beyond current hardware size limitations.
While encouraging heuristic approaches such as the quantum-classical sampling algorithm described here continue to be developed across the optimization community, we can see there is much optimism to be found in anticipating utility, and even advantage, before fault-tolerance in quantum optimization.
@@ Appendix A Boltzmann sampling
In preparation for using QeMCMC in the heuristic setting implemented in this paper, we conduct small-scale Boltzmann sampling experiments for various temperatures, as were performed in Ref. [41]. We do this using a small, five-node MIS problem instance. Owing to the modest dimensionality of this system, the complete distribution over computational basis states can be readily visualized. In Figure 5, we see the theoretical Boltzmann distribution for this problem instance. Here, the Ising-like energies are computed by enumerating all $2^{n}$ bitstrings $\{s_{i}\}$ and evaluating their energies $E_{i}=E(s_{i})$ under the MIS Hamiltonian. The corresponding probabilities are then obtained from the Boltzmann distribution defined in Eq. (1), yielding the exact classical distribution at temperature $T$. In Figure 6(a), we demonstrate the iterative process of converging to the desired Boltzmann distributions at $T=0.1$, where $k_{B}=1$ on IBM Quantum superconducting hardware. When Ref. [41] was originally published, the authors were limited from running these experiments end-to-end, within one execution window. This was due to limitations of running iterative experiments on the available quantum hardware, which is explained fully in the associated work [41]. However, this execution style is now well-supported with Qiskit Runtime Session mode. In the work presented here, we use Session mode alongside parameter streaming capabilities, which provide a tight classical feedback loop for determining circuit parameters based on quantum experimental results, within the Session window. The results shown in Figure 6(a) and Figure 6(b) were obtained in this way, which allows for efficient job batching and reduced queuing latency.
@@
Next, we apply the MH acceptance probability as defined in Eq. (4), to construct a transition matrix, $P$, from the proposal matrix, $Q$, equivalent to $P$ in Eq. (3). The resulting transition matrix is stochastic, and fully describes our Markov chain, incorporating both the proposal distribution and the MH acceptance rule.
Our transition matrix, $P$, is then used to generate moves in the Markov chain, and over many iterations, or steps, we can construct plots like that shown in Figure 6(a) and in Figure 6(b). From this, we can check that our stationary distribution is an eigenstate of the transition matrix, $Q$.
We note that there is a general assumption within the community that high-performance computing (HPC) resources are required for implementing QeMCMC at non-trivial scales. In our work, HPC resources were used for a number of tasks both in experiment preparation and post-processing results, such as with parameter optimization and classical simulation experiments of the larger quantum circuits used herein. However, generally, we avoided excessive use of HPC resources during the algorithm execution window by pre-emptively compiling circuits and using known QAOA best practices to scale our experiments. We also acknowledge that QeMCMC algorithms generally rely on a fast exchange of information between sampling from the quantum computer and a classical evaluation procedure to accept or reject a proposal. For the scales of problems investigated here, up to $117$ qubits, the use of remote APIs, such as the Qiskit Runtime Session mode mentioned previously, were sufficient for our use case. However, we anticipate the classical computational demand of this algorithm to increase as we look to solve larger problem instances.
@@ Appendix B Maximum independent set mapping
again up to an additive constant.
Here, the first term is a constant energy offset, while the remaining terms define the Hamiltonian encoding the constraints of the maximum independent set problem. These terms consist of single-qubit contributions proportional to $Z_{i}$ and pairwise interaction terms of the form $Z_{i}Z_{j}$. Upon promoting the classical spin variables to Pauli-$Z$ operators, the independent set objective and constraints take the form of a classical Ising Hamiltonian,
@@
where the coefficients $h_{i}$ and $J_{ij}$ arise directly from the linear and quadratic operator terms in Eqs. (20) and (21), up to a similar additive constant.
Finally, while the MIS problem is naturally posed as a maximization of the objective $\sum_{i}x_{i}$, we adopt the standard physics convention of defining an energy function as the negative of this objective (with additional penalty terms), thereby converting the problem into an equivalent energy minimization task whose ground states correspond to optimal solutions.
@@ Appendix D Tuning the parallel tempering experimental setup
In this Appendix, we explain how the parameters used in the parallel tempering setup were selected. As a reminder, parallel tempering was initially developed to improve MCMC sampling of Gibbs distributions [52, 29, 34], but it is also commonly adapted as an optimization tool to find low‑energy or ground‑state solutions, where we expect the target ground state energy to coincide with the expected value of the Gibbs free energy of a low temperature replica of the parallel tempering experiment [16, 58]. Accordingly, we use several standard design considerations from the parallel tempering literature [16, 52, 29, 34, 30, 40, 42, 47, 7] to guide our experiments.
Firstly, determining the number of replicas and their corresponding temperatures is key in ensuring algorithmic performance [16]. In parallel tempering, the maximum temperature $T_{\text{high}}$ should be chosen so that replicas can escape local energy minima, while the number of replicas must be sufficient to allow swaps to occur reliably between adjacent temperature levels. In our setup, we determined the temperature range by first computing the maximum temperature using an inverted Boltzmann factor,

### v2

@@ V Discussion & Conclusions
In this manuscript, we explored a quantum optimization algorithm based on a QeMCMC sampling routine to solve non-trivial combinatorial optimization problem instances. We have adapted the QeMCMC workflow to include a warm-started quantum optimization circuit and an intentional bias of our proposal distribution towards high-quality solutions. We further combined this with a parallel tempering approach, which is commonly used to find ground states of many-body systems, to explore the available cost landscape more efficiently.
This heuristic quantum-classical algorithm has been tested against problem instances with up to 117 decision variables, corresponding to 117 qubits on real quantum hardware. While our method does not offer provable convergence guarantees, we recover the global optimum for all considered problem instances with 17, 39, 52, 80 and 117 variables. The results provide early evidence of a scaling advantage over comparable classical MCMC approaches for the chosen problem instances. Notably, the quantum hardware also outperforms MPS simulations under the considered settings.
While these results provide a promising start, several open questions remain. More experiments across larger sets of instances and additional problem classes are needed to further strengthen our conclusions. Alternative circuit constructions may further improve the performance, and approaches for scaling beyond hardware limitations, such as problem decomposition or block sampling, should be investigated.
To conclude, heuristic quantum‑enhanced sampling approaches such as using the one explored here highlight growing optimism that practical utility, and even advantage, in quantum optimization may be achievable before full fault tolerance.
@@ Appendix A Boltzmann sampling
In preparation for using QeMCMC in the heuristic setting implemented in this paper, we conduct small-scale Boltzmann sampling experiments for various temperatures, as were performed in Ref. [42]. We do this using a small, five-node MIS problem instance. Owing to the modest dimensionality of this system, the complete distribution over computational basis states can be readily visualized. In Figure 5, we see the theoretical Boltzmann distribution for this problem instance.
Here, the Ising-like energies are computed by enumerating all $2^{n}$ bitstrings $\{s_{i}\}$ and evaluating their energies $E_{i}=E(s_{i})$ under the MIS Hamiltonian. The corresponding probabilities are then obtained from the Boltzmann distribution defined in Eq. (1), yielding the exact classical distribution at temperature $T$.
In Figure 6(a), we demonstrate the iterative process of converging to the desired Boltzmann distributions at $T=0.1$, where $k_{B}=1$ on IBM Quantum superconducting hardware. When Ref. [42] was originally published, the authors were limited from running these experiments end-to-end, within one execution window. This was due to limitations of running iterative experiments on the available quantum hardware, which is explained fully in the associated work [42]. However, this execution style is now well-supported with Qiskit Runtime Session mode. In the work presented here, we use Session mode alongside parameter streaming capabilities, which provide a tight classical feedback loop for determining circuit parameters based on quantum experimental results, within the Session window. The results shown in Figure 6(a) and Figure 6(b) were obtained in this way, which allows for efficient job batching and reduced queuing latency.
@@
Next, we apply the MH acceptance probability as defined in Eq. (4), to construct a transition matrix, $P$, from the proposal matrix, $Q$, equivalent to $P$ in Eq. (3). The resulting transition matrix is stochastic, and fully describes our Markov chain, incorporating both the proposal distribution and the MH acceptance rule. Our transition matrix, $P$, is then used to generate moves in the Markov chain, and over many iterations, or steps, we can construct plots like that shown in Figure 6(a) and in Figure 6(b). From this, we can check that our stationary distribution is an eigenstate of the transition matrix, $Q$.
We note that there is a general assumption within the community that HPC resources are required for implementing QeMCMC at non-trivial scales. In our work, HPC resources were used for a number of tasks both in experiment preparation and post-processing results, such as with parameter optimization and classical simulations of our largest implemented quantum circuits.
However, generally, we avoided excessive use of HPC resources during the algorithm execution window by pre-emptively compiling circuits and using known QAOA best practices to scale our experiments. We also acknowledge that QeMCMC algorithms generally rely on a fast exchange of information between sampling from the quantum computer and a classical evaluation procedure to accept or reject a proposal. For the scales of problems investigated here, up to $117$ qubits, the use of remote APIs, such as the Qiskit Runtime Session mode mentioned previously, were sufficient for our use case. However, we anticipate the classical computational demand of this algorithm to increase as we look to solve larger problem instances.
@@ Appendix B Maximum independent set mapping
again up to an additive constant. Here, the first term is a constant energy offset, while the remaining terms define the Hamiltonian encoding the constraints of the maximum independent set problem. These terms consist of single-qubit contributions proportional to $Z_{i}$ and pairwise interaction terms of the form $Z_{i}Z_{j}$. Upon promoting the classical spin variables to Pauli-$Z$ operators, the independent set objective and constraints take the form of a classical Ising Hamiltonian,
@@
where the coefficients $h_{i}$ and $J_{ij}$ arise directly from the linear and quadratic operator terms in Eqs. (20) and (21), up to a similar additive constant. Finally, while the MIS problem is naturally posed as a maximization of the objective $\sum_{i}x_{i}$, we adopt the standard physics convention of defining an energy function as the negative of this objective (with additional penalty terms), thereby converting the problem into an equivalent energy minimization task whose ground states correspond to optimal solutions.
@@ Appendix D Tuning the parallel tempering experimental setup
In this Appendix, we explain how the parameters used in the parallel tempering setup were selected. As a reminder, parallel tempering was initially developed to improve MCMC sampling of Gibbs distributions [55, 29, 35], but it is also commonly adapted as an optimization tool to find low‑energy or ground‑state solutions, where we expect the target ground state energy to coincide with the expected value of the Gibbs free energy of a low temperature replica of the parallel tempering experiment [16, 61].
Accordingly, we use several standard design considerations from the parallel tempering literature [16, 55, 29, 35, 30, 41, 43, 48, 7] to guide our experiments. Firstly, determining the number of replicas and their corresponding temperatures is key in ensuring algorithmic performance [16]. In parallel tempering, the maximum temperature $T_{\text{high}}$ should be chosen so that replicas can escape local energy minima. The number of replicas must be sufficient to allow swaps to occur reliably between adjacent temperatures. In our setup, we determined the temperature range by first computing the maximum temperature using an inverted Boltzmann factor,